OpenStudy (anonymous):

(3x-6) (x+8) find the Product ...

OpenStudy (anonymous):

do you know how to use FOIL?

OpenStudy (anonymous):

no.. sorry

OpenStudy (anonymous):

Okay, it stands for First, Inner, Outer, Last. (3x-6) (x+8) Multiply the first terms in the binomial, tell me what you get.

OpenStudy (anonymous):

The first terms in each binomial are 3x and x.

OpenStudy (anonymous):

so.. 3x^2?

OpenStudy (anonymous):

Yes, so we've done the F in FOIL

OpenStudy (anonymous):

next..

OpenStudy (anonymous):

Now we have OIL, Outer is next. The terms that are farthest from each other. Do you see what they are?

OpenStudy (anonymous):

24x?

OpenStudy (anonymous):

Yes! :D

OpenStudy (anonymous):

So we have done FO in FOIL, next is Inner. The terms that are closest to each other and not the first or last term. derp.

OpenStudy (anonymous):

6x?

OpenStudy (anonymous):

Think about it a little more... what's the first binomial?

OpenStudy (anonymous):

-6x?

OpenStudy (anonymous):

:D You got it, now the last. The last terms in each binomial.

OpenStudy (anonymous):

-48?

OpenStudy (anonymous):

Yup! Now put what you have got together.

OpenStudy (anonymous):

3x^2+24x -6-48

OpenStudy (anonymous):

You mistyped a bit, but yeah, 3x^2 - 6x + 24x - 48. Now what do you think you have to do?

OpenStudy (anonymous):

dont know,,,

OpenStudy (anonymous):

it should be quadratic..

OpenStudy (anonymous):

The last step in almost every single equation in Mathematics. o.o SIMPLIFY!!! WOOOO

OpenStudy (anonymous):

You have to terms of the first degree, which means you can combine them. :D

OpenStudy (anonymous):

two*. derp.

OpenStudy (anonymous):

..

OpenStudy (anonymous):

3x^2 - 6x + 24x - 48 3x^2 + 18x - 48

OpenStudy (anonymous):

See it now?

OpenStudy (anonymous):

ahh.. thanks :)

OpenStudy (anonymous):

No problem buddy. :) Just remember, FOIL. F = (De)(Rp) O = (De)(rP) I = (dE)(Rp) L = (dE)(rP)